What is the equation of the normal to the curve y = sin x at (0, 0)?

(a)x =0 (b) y=0 (c)x+y =0 (d)x-y=0

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Given that, y = sin x

Hence, dy/dx = cos x

Thus, the slope of the normal = (-1/cos x)@x =0= -1

Therefore, the equation of the normal is y-0 = -1(x-0) or x+y=0

the correct solution is option C.
Thanks